What is the activity coefficient for each ion at the given ionic strength? Click here for an activity coefficient table.What is the activity coefficient for each ion at t
Answer #1

This problem involves looking at a table given the species, their charge and their ionic strenght. You can find the table of the activity coefficient on the web by just typing "activity coefficient table".

We can sketch the whole problem in one single table for more clarity:

Species Charge Ionic strenght Activity coefficient
HPO42- -2 0.01M 0.660
Y3+ +3 0.005M 0.54
Pr3+ +3 0.1M 0.18
(CH3CH2)3NH+ +1 0.01M 0.904
Answer #2

You can find the answer by applying the next equation:

log\gamma = \frac{-0.51*z^{2}*\sqrt{\mu }}{1+\left ( \alpha \sqrt{\mu }/305 \right )}

alpha is the size of ion (from table)

Miu is the ionic strenght (given)

z is the charge (given)

Applying this equation for every compound I get.

I + (α 305 Miu sqrt alfa log activity co activity co 400 -0.18034783 0.660164511 miu -2 40.01 0.1 hpo4-2 9 0.005 0.07071068 9